OpenStudy (anonymous):

for a population of 800000 subway riders, the numbers of subway trips taken per rider last january are approximately normally distributed with a mean of 56 trips and a standard deviation of 13 trips. approximately how many of the riders took between 30 and 43 trips last january???

OpenStudy (anonymous):

I hate statistics

OpenStudy (anonymous):

me too studyn for my gre

OpenStudy (amistre64):

ahhh, zscores, gotta love em

OpenStudy (amistre64):

what do you have for computing, table or calculator?

OpenStudy (anonymous):

table..and calc on my laptop

OpenStudy (amistre64):

\[Z=\frac{x-\mu}{\sigma}\]

OpenStudy (amistre64):

we wanna cram the non standard stuff into the normal curve

OpenStudy (anonymous):

I neva seen that z formula before

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the zscore allows us to compare different amounts as tho they were equals

OpenStudy (anonymous):

ok..

OpenStudy (amistre64):

so lets start cramming it in :) mean = 56 30 - 56 = |-26| 43 - 56 = |-13| and sd = 13, were in luck, your numbers are divisable by 13 so its fairly simple

OpenStudy (amistre64):

26/13 = 2, we are 2 standard deviations from the mean 13/13 = 1, we are 1 standard deviation from the mean the area between 1 and 2 is what we want to find

OpenStudy (amistre64):

do you know the "empirical rule"?

OpenStudy (anonymous):

ok...so the answers given are 60000 110000 160000 210000 270000

OpenStudy (anonymous):

no explain empirical rule..

OpenStudy (amistre64):

i got no idea what the answer is yet, I am going thru it along with you :)

OpenStudy (amistre64):

the empirical rule is a rule of thumb for the area/probability under the curve for deviations from the mean of 1,2, and 3

OpenStudy (amistre64):

1 = 68 2 = 95 3 = 99 but, we need only half of these since our sds are just on one side

OpenStudy (amistre64):

2 - 1 will give us the approximate area 95/2 - 68/2 = N 95 -68 ---- 27 /2 = 13.5 right?

OpenStudy (amistre64):

if so then our answer is: 800,000 * .135

OpenStudy (anonymous):

.135 * 800000= 108000 looks close to 110000 thanks..

OpenStudy (amistre64):

looks good to me, and it makes sense as well :) hope its right ;)

OpenStudy (anonymous):

so situations like these i dont necessarily need to draw a bell curve diagam??

OpenStudy (amistre64):

draw the curve, its good practice for later

OpenStudy (amistre64):

and understand the z score / sds

OpenStudy (amistre64):

zscore relates to standard deviations as: 1 zscore: 1 sd in this problem we are given an sd of 13 so thats a ration of; 1:13, or simply put 1/13 when the mean is at 0 we can equate the distance of any given data point from the mean by using this ratio. take for instance the data given here: mean = 56 and we have a distance of 30 away from it; 30 <---->56 -56 -56 ---- ----- -26 <----> 0 26 we simply shift it all to look at it at a mean of 0, and an sd of 1, so use our ratio now of 1/13 26*1z/13 = 26z/13 = 2z, we have a zscore of 2, or simply 2 sds away from the mean

OpenStudy (amistre64):

not a "distance of 30" but rather we need to find the distance between 56 and 30 ....

OpenStudy (anonymous):

ahhhh thankkkkks